# Revisit, Extend, and Enhance Hessian-Free Influence Functions

### v1

@@ 5 Noisy Label Correction for Vision Datasets
In this section, we demonstrate the effectiveness of our IP Ensemble in identifying samples with noisy labels on vision datasets. Specifically, we choose three benchmark datasets CIFAR-10N [46], CIFAR-100N [46], and Animal-10N datasets [47] in the noisy label learning area, all of which are derivatives of the original CIFAR-10 [48], CIFAR-100 [48], and Animal-10 [49], but incorporate label inaccuracies from crowd-sourced annotation. CIFAR-10N encompasses three distinct noise settings: aggregate, random, and worst, denoted as “-a", “-r", and “-w", respectively. In the aggregate setting, labels are derived via majority voting among three annotators, with ties being resolved randomly. The random setting adopts the label provided by the first annotator, while the worst setting selects the label from the least reliable annotator.
@@
For the competitive methods, we choose the following influence function-based methods. TracIn [16] replaces the Hessian matrix with the identity matrix; LiSSA [10] and EKFAC [13] employ implicit Hessian-vector products and Kronecker-Factored curvature to efficiently approximate the inverse of the Hessian matrix; DataInf [3] swaps the order of the matrix multiple for obtaining a closed-form expression; Self-TracIn [18] and Self-LiSSA [17] are the self-expression versions of TracIn and LiSSA, where $\nabla v^{\textrm{util}}$ is replaced with $\nabla_{\hat{\theta}}\ell(z_{i};\hat{\theta})$, leading their self-influence to $\ell(z_{i};\hat{\theta})^{\top}\nabla_{\hat{\theta}}\ell(z_{i};\hat{\theta})$ or $\ell(z_{i};\hat{\theta})^{\top}\mathbf{H}^{-1}_{\hat{\theta}}\nabla_{\hat{\theta}}\ell(z_{i};\hat{\theta})$. Our IP closely resembles TracIn, with the additional application of dropout with $\mathcal{U}(0,0.01)$ on the model parameters. Furthermore, our IP Ensemble constitutes an ensemble version of IP with an ensemble size of 5.
To reduce randomness, we train a ResNet-34 network [50] once on each dataset to establish a baseline model. Subsequently, based on the same model, we employ the aforementioned influence function-based methods to identify 5% of detrimental samples. Following this, we conduct five retraining iterations of the ResNet-34 network, each time removing the identified detrimental samples from the training set. We report the average accuracy and standard deviation of the above influence function-based methods across these five retraining iterations in Table 1. In general, these influence function-based methods are effective in identifying detrimental samples. Upon retraining the ResNet-34 model without these identified samples, nearly every result obtained by these methods outperforms the vanilla ResNet-34 trained on the entire dataset, except for EKFAC on CIFAR-10N-w. Notably, our IP Ensemble consistently outperform other baseline methods across various noise conditions and datasets. Particularly noteworthy is the performance of our IP Ensemble in the most challenging scenario, CIFAR-100N, achieving the highest recorded accuracy of 62.25% on the test set, surpassing the vanilla cross-entropy accuracy of 56.41%. Additionally, compared to TracIn, the non-ensemble version of IP, IP Ensemble delivers further performance gains, highlighting the benefits of the ensemble strategy for enhancing model generalization. Moreover, the average accuracy of the IP Ensemble on the test set reaches its peak at 82.93%, surpassing both the vanilla cross-entropy performance of 80.90% and the second-best accuracy of Self-TracIn at 82.45%. More experimental results and analysis on different percentage of removed samples, parameter analysis on different dropout rates and different ensemble size can be found in Appendix C.2 and C.3, respectively.
@@
In addition to the algorithmic performance mentioned above, we also present the running times of various influence function-based methods for identifying detrimental samples. Despite some baselines having linear time complexity, there is significant divergence in their real execution times. Given that TracIn and our IP Ensemble exhibit exceptional speed and similar execution times, we consider them as the baseline and compute the speed improvement factors over other baseline methods, as depicted in Figure 4. With the exception of Self-TracIn, our IP runs over 100 times faster than LiSSA, EKFAC, DataInf, and Self-LiSSA. Notably, on Animal-10N, IP is over 800 times faster than EKFAC. It’s worth noting that Self-TracIn and Self-LiSSA are slower than their standard versions due to the fixed $\nabla v^{\textrm{util}}$ when calculating the influence of each sample. The simplicity and efficiency of IP make it a promising tool for analyzing sample influence on deep models. The time complexities and execution time of these methods can be found in Appendix C.4.
@@
Figure 5: Accuracy and fair score of different influence function-based methods on fine-tuning RTE, COLA, and QNLI datasets. The X-axis denotes accuracy, and the Y-axis for fair score is inverted. The brown crossing denotes the performance of using all the samples for fine-tuning the RoBERTa model as the baseline model. Building upon this, we plot a horizontal and a vertical line in each figure and divide the space by fairness and utility results into four regions. The green area in the top right corner signifies a model that is both fairer and more accurate compared to the baseline model.
@@ 6 Data Curation Towards Fine-Tuning of Fairer NLP Models
In this section, we further demonstrate the efficacy of our IP method in gauging the impact of individual samples on fairness within the realm of curating suitable data samples for fine-tuning Natural Language Processing (NLP) models. Beyond mere utility, fairness has emerged as an indispensable attribute for machine learning models to mitigate inadvertent discrimination.
@@
In this experiment, we employ three datasets—RTE, COLA, and QLIN—from the GLUE repository [51] to fine-tune the widely-used language model pre-training model RoBERTa [52]. Our focus is on group fairness, necessitating that machine learning models treat samples within various predefined subgroups comparably. To assess fairness, we adopt the methodology outlined in [53], which involves perturbing the demographic information within each sample and scrutinizing whether the model yields identical predictions for the original sample $x$ and its corresponding perturbed counterpart $\tilde{x}$. This evaluation metric "fair score" is defined as $|\mathcal{C}(x)-\mathcal{C}(\tilde{x})|$ over all the samples in the test set, normalized by the size of the test set, where $\mathcal{C}(\cdot)$ is the model predictor. It is worth to note that the fair score is a negative metric, indicating that smaller values are preferable.
Utility and fairness serve as distinct perspectives for evaluating the performance of a model. While a constant predictor that consistently outputs the same prediction, regardless of input, can achieve perfect fairness, it lacks practical flexibility. Thus, in our fine-tuning experiments, we consider both utility and fairness. Employing the same influence function-based methods and our IP Ensemble as in the previous section, we conduct a comparative analysis. For each method, we calculate the influence on both utility and fairness within the fine-tuning set. We then identify and remove the 5% most detrimental samples in terms of utility and fairness before fine-tuning the RoBERTa model to optimize performance across both accuracy and fairness metrics.
@@
The results of this experiment are presented in Figure 5, where the Y-axis for fairness is inverted. The brown crossing denotes the performance of using all the samples for model fine-tuning as the base model. Building upon this, we plot a horizontal and a vertical line in each figure and divide the space by fairness and utility results into four regions. The green area in the top right corner signifies a model that is both fairer and more accurate compared to the baseline model. Results within this green region can be considered Pareto improvements, enhancing both utility and fairness simultaneously. It is evident most results are located in the green area, indicating the existence of detrimental samples, and not all the samples are helpful to the model performance. This also implies that influence function methods are effective to identifying the detrimental samples, even for non-convex deep models. However, it is important to note that some competitive methods yield a trade-off or even Pareto deterioration. For instance, LiSSA demonstrates a better fair score but worse accuracy compared to the base model on RTE; conversely, it exhibits better accuracy but worse fairness on QLIN. EKFAC shows similar performance on COLA and QLIN. Self-LiSSA demonstrates Pareto deterioration on COLA. We hypothesize that the approximation of the inverse of the Hessian matrix may suffer from large errors, leading to heavily divergent influence estimations from the true values. Our IP Ensemble consistently achieves Pareto improvements across all three datasets, often yielding the best results compared to other influence function-based methods.
@@
7 Defense Against Adaptive Evasion Adversaries
In this experiment, we utilize a Logistic Regression model with three datasets: Bank [56], CelebA [57], and JigsawToxicity [58]. Following the protocol in Section 3, we consider a white-box adversary [44] to craft adversarial samples. For each sample $(x_{j},y_{j})$ in the validation set $\mathcal{V}$, we perturb it by changing the feature $x^{\prime}_{j}=x_{j}-\gamma\frac{\hat{\theta}^{\top}x_{j}+b}{\hat{\theta}^{\top}\hat{\theta}}\hat{\theta}$ and keeping $y_{j}$ unchanged. The attacker perturbs between 5% and 25% of the test set samples at random. By quantifying the impact of samples on model robustness, we trim 5% detrimental samples in the training set through influence functions. The boxplot depicted in Figure 6 demonstrates the performance variations across various influence function-based methods over 10 distinct attacks. In general, influence function-based methods are highly effective against adaptive evasion attackers, showcasing superior resilience particularly post-attack, notably evident in scenarios involving Bank and JigsawToxicity. Among these influence function-based methods, our IP Ensemble demonstrates competitive efficacy with the best or the second best on these three datasets.
In addition to the above removing strategy, we continue to explore relabeling and reweighting strategies to tackle detrimental samples. The relabeling strategy [20] changes the identified detrimental samples from their original classes into another class, since the datasets we used here are binary classes, we directly flip their labels. The reweighting strategy [18] takes the influence score of each sample as the exponential weight with a softmax normalization, and then trains a model with weighted gradient. We report the defense performance of various influence function-based methods and our IP Ensemble under the relabeling and reweighting strategies in Table 2.
Under the relabeling strategy, DataInf achieves the best results on Bank and CelebA with 87.46% and 77.36% accuracy, respectively. However, LiSSA is not stable as other methods on CelebA with only 68.68%, which is even worse than the performance of post attack. Our IP ensemble method continues to achieve competitive performance, with accuracy scores of 87.45%, 77.30%, and 70.51% on three datasets with the second-best performance among all influence functions-based methods. For the reweighting strategy, the influence function-based methods are still effective on Bank and JigsawToxicity, which outperform the post-attack by a large margin, but achieve similar performance with post attack on CelebA. Notably, when considering the average accuracy of both the relabeling and reweighting strategies, our IP Ensemble method performs the best.
@@ 8 Conclusion
In this paper, we revisited, extended, and enhanced the TracIn method into the Inner Product (IP) formulation. By proving the exhibition of order-consistency in several scenarios between IP and Influence function, we demonstrated how substituting the inverse of the Hessian matrix with an identity matrix offers a practical and computationally efficient solution to estimating sample influence. Based on that, we extended our IP to measure the sample influence on fairness and robustness. Continually, we enhanced the generalization of IP by introducing an ensemble strategy. We verified the correctness of IP on synthetic datasets and extensive evaluations on noisy label detection, data curation for fair NLP model fine-tuning, and defense against adaptive adversarial attacks. Overall, our IP Ensemble highlighted the potential of simple, yet powerful, approximations in influence estimation. By broadening the scope of TracIn and improving its robustness and fairness applications, we provided valuable tools for data-centric learning and model interpretation, offering practical solutions for contemporary challenges in deep learning.
@@ Appendix A Rationality of Inner Product
  • (I, I). For a case where both $z_{i}$ and $z_{i^{\prime}}$ have gradients in Region I, $\mathcal{I}^{\textrm{util}}(-z_{i})\gtrless\mathcal{I}^{\textrm{util}}(-z_{i^{\prime}})$ implies $\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i^{\prime}})\gtrless\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i})$, which is the opposite of order-consistency. While these samples are jointly recognized by $\mathcal{I}^{\textrm{util}}$ and $\mathcal{I}^{\textrm{util}}_{\textrm{IP}}$ as detrimental, there is little practical difference.
  • (I, II). For a case where $z_{i}$ has gradients in Region I and $z_{i^{\prime}}$ has gradients in Region II, the order-consistency partially holds, depending on which Sub-Region the gradients of $z_{i^{\prime}}$ is in. For Sub-Region $\mathrm{II^{b}}$ and $\mathrm{II^{c}}$, we have $\mathcal{I}^{\textrm{util}}(-z_{i})>\mathcal{I}^{\textrm{util}}(-z_{i^{\prime}})$ and $\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i})>\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i^{\prime}})$, and the order-consistency holds. For Sub-Region $\mathrm{II^{a}}$, however, $\mathcal{I}^{\textrm{util}}(-z_{i})\gtrless\mathcal{I}^{\textrm{util}}(-z_{i^{\prime}})$ and $\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i})>\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i^{\prime}})$, thus the order-consistency partially holds. While samples in Sub-Region $\mathrm{II^{a}}$ are jointly recognized by $\mathcal{I}^{\textrm{util}}$ and $\mathcal{I}^{\textrm{util}}_{\textrm{IP}}$ as detrimental, there is little practical difference.
  • (I, III). For a case where $z_{i}$ has gradients in Region I and $z_{i^{\prime}}$ has gradients in Region III, $\mathcal{I}^{\textrm{util}}(-z_{i})>\mathcal{I}^{\textrm{util}}(-z_{i^{\prime}})$ and $\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i})>\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i^{\prime}})$, therefore the order-consistency holds.
  • (II, IV). For a case where $z_{i}$ has gradients in Region II and $z_{i^{\prime}}$ has gradients in Region IV, the order-consistency partially holds.
  • (II, II). For a case where both $z_{i}$ and $z_{i^{\prime}}$ have gradients in Region II, $\mathcal{I}^{\textrm{util}}(-z_{i})\gtrless\mathcal{I}^{\textrm{util}}(-z_{i^{\prime}})$ implies $\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i})\gtrless\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i^{\prime}})$, the order-consistency holds.

### v2

@@ 5 Noisy Label Correction for Vision Datasets
In this section, we demonstrate the effectiveness of our IP Ensemble in identifying detrimental samples on noisy vision datasets. Specifically, we choose three benchmark datasets CIFAR-10N [60], CIFAR-100N [60], and Animal-10N datasets [50] in the noisy label learning area. CIFAR-10N encompasses three distinct noise settings: aggregate, random, and worst, denoted as “-a,” “-r,” and “-w,” respectively. “a” means that labels are derived via majority voting among three annotators, with ties being resolved randomly, “r” adopts the label provided by the first annotator, while “w” selects the label from the least reliable annotator.
For competitive methods, we choose the following influence function-based methods. TracIn [46] replaces the Hessian matrix with the identity matrix and considers checkpoints during the training process; LiSSA [26] and EKFAC [19] employ implicit Hessian-vector products and Kronecker-Factored curvature to efficiently approximate the inverse of the Hessian matrix; DataInf [31] swaps the order of the matrix multiple for obtaining a closed-form expression; Self-TracIn [52] and Self-LiSSA [5] are the self-expression versions of TracIn and LiSSA, where $\nabla v^{\textrm{util}}$ is replaced with $\nabla_{\hat{\theta}}\ell(z_{i};\hat{\theta})$ and only the last checkpoint, i.e., the converged model parameters, is used in Self-TracIn. GEX [25] utilizes ensemble methods based on checkpoints from extra stochastic gradient descent on the converged model. TDA [3] focuses checkpoints during the training process, ensembles the influence via EKFAC [19]. Our IP Ensemble method constitutes an ensemble version of IP with $\mathcal{U}(0,0.01)$ dropout applied on model parameters and an ensemble size of 5.
To reduce randomness, we train a ResNet-34 network [22] once on each dataset to establish a baseline model. Subsequently, based on the same model, we employ the aforementioned influence function-based methods to identify 5% of detrimental samples. Following this, we conduct five retraining iterations of the ResNet-34 network, each time removing the identified detrimental samples from the training set. We report the average accuracy and standard deviation of the above influence function-based methods across these five retrainings in Table 1. In general, these influence function-based methods are effective in identifying detrimental samples. Upon retraining the ResNet-34 model without these identified samples, nearly every result obtained by these methods outperforms the vanilla ResNet-34 trained on the entire dataset, except for EKFAC on CIFAR-10N-w. Notably, our IP Ensemble consistently outperforms other baseline methods across various noise conditions and datasets. Particularly noteworthy is the performance of our IP Ensemble in the most challenging scenario, CIFAR-100N, achieving the highest recorded accuracy of 62.25% on the test set, surpassing the vanilla cross-entropy accuracy of 56.41%. Moreover, TDA and TracIn only achieve moderate performance because they introduce multiple checkpoints throughout the training process. These approaches might dilute the effectiveness of identifying truly detrimental samples for the converged model, as the influence of each checkpoint may vary. Additionally, compared to IP, IP Ensemble delivers further performance gains, highlighting the benefits of the ensemble strategy for enhancing model generalization. Moreover, the average accuracy of the IP Ensemble on the test set reaches its peak at 82.93%, surpassing both the vanilla cross-entropy performance of 80.90% and the second-best accuracy of Self-TracIn at 82.45%. More experimental results and analysis on different percentages of removed samples, parameter analysis on different dropout rates, ensemble sizes, and base model architectures can be found in Appendix C.2C.4, respectively.
@@
In addition, we also present the running times of these influence function-based methods. Despite some baselines having linear time complexity, there is significant divergence in their real execution times. Given that our IP exhibits exceptional speed and similar execution times, we consider them as the baseline and compute the speed improvement factors over other baseline methods, as depicted in Figure 4. For ensemble methods including TracIN, TDA, GEX, and our IP Ensemble, parallel computation can be applied to accelerate the running time if enough resources are allowed; if calculated serially, the time grows linearly with the ensemble size. We do not report their running time in Figure 4. With the exception of Self-TracIn, our IP runs over 100 times faster than LiSSA, EKFAC, DataInf, and Self-LiSSA. Notably, on Animal-10N, IP is over 800 times faster than EKFAC. It is worth noting that Self-TracIn and Self-LiSSA are slower than their standard versions due to the fixed $\nabla v^{\textrm{util}}$ when calculating the influence of each sample. The simplicity and efficiency of IP make it a promising tool for analyzing sample influence on deep models. The time complexities and execution time of these methods can be found in Appendix C.5.
@@
Figure 5: Accuracy and fair score of different influence function-based methods on fine-tuning RTE, COLA, and QNLI datasets. The X-axis denotes accuracy, and the Y-axis for fair score is inverted. The brown crossing denotes the performance of using all the samples for fine-tuning the RoBERTa model as the baseline model. Building upon this, we plot a horizontal and a vertical line in each figure and divide the space by fairness and utility results into four regions. The green area in the top right corner signifies a model that is both fairer and more accurate compared to the baselines.
@@ 6 Data Curation Towards Fine-Tuning of Fairer NLP Models
In this section, we further demonstrate the efficacy of our IP method in gauging the impact of individual samples on fairness within the realm of curating suitable data samples for fine-tuning language models. Beyond mere utility, fairness has emerged as an indispensable attribute for machine learning models to mitigate inadvertent discrimination.
@@
In this experiment, we employ three datasets—RTE, COLA, and QLIN—from the GLUE repository [56] to fine-tune the widely-used language model RoBERTa [35]. Our focus is on group fairness, necessitating that machine learning models treat samples within various predefined subgroups comparably. To assess fairness, we adopt the methodology outlined in [47], which involves perturbing the demographic information within each sample and scrutinizing whether the model yields identical predictions for the original sample $x$ and its corresponding perturbed counterpart $\tilde{x}$. This evaluation metric "fair score" is defined as $|\mathcal{C}(x)-\mathcal{C}(\tilde{x})|$ over all the samples in the test set, normalized by the size of the test set, where $\mathcal{C}(\cdot)$ is the model predictor. Note that the fair score is a negative metric, so smaller values are preferable. Utility and fairness serve as distinct perspectives for evaluating the performance of a model. Thus, in our fine-tuning experiments, we consider both utility and fairness. Employing the same influence function-based methods and our IP Ensemble as in the previous section, we conduct a comparative analysis. For each method, we calculate the influence on both utility and fairness within the fine-tuning set. Subsequently, We identify and remove the 5% most detrimental samples in terms of utility and fairness before fine-tuning the RoBERTa model to optimize performance across both accuracy and fairness metrics.
@@
The results of this experiment are presented in Figure 5, where the Y-axis for fairness is inverted. The brown crossing denotes the performance of using all the samples for model fine-tuning as the base model. Building upon this, we plot a horizontal and a vertical line in each figure and divide the space by fairness and utility results into four regions. The green area in the top right corner signifies a model that is both fairer and more accurate compared to the baseline model. Results within this green region can be considered Pareto improvements, enhancing both utility and fairness simultaneously. It is evident most results are located in the green area, indicating the existence of detrimental samples, and not all the samples are helpful to the model performance. This also implies that influence function methods are effective to identifying the detrimental samples, even for non-convex deep models. However, it is important to note that some competitive methods yield a trade-off or even Pareto deterioration. For instance, LiSSA demonstrates a better fair score but worse accuracy compared to the base model on RTE; conversely, it exhibits better accuracy but worse fairness on QLIN. EKFAC shows similar performance on COLA and QLIN. Self-LiSSA demonstrates Pareto deterioration on COLA. We hypothesize that the approximation of the inverse of the Hessian matrix may suffer from large errors, leading to heavily divergent influence estimations from the true values. Our IP Ensemble consistently achieves Pareto improvements across all three datasets, yielding the best results compared to other influence-based methods.
7 Defense Against Adaptive Evasion Adversaries
In this experiment, we utilize a Logistic Regression model with three datasets: Bank [40], CelebA [37], and JigsawToxicity [41]. Following the protocol in Section 3, we consider a white-box adversary [39] to craft adversarial samples. For each sample $(x_{j},y_{j})$ in the validation set $\mathcal{V}$, we perturb it by changing the feature $x^{\prime}_{j}=x_{j}-\gamma\frac{\hat{\theta}^{\top}x_{j}+b}{\hat{\theta}^{\top}\hat{\theta}}\hat{\theta}$ and keeping $y_{j}$ unchanged. The attacker perturbs between 5% and 25% of the test set samples at random. By quantifying the impact of samples on model robustness, we trim 5% detrimental samples in the training set through influence functions. The boxplot depicted in Figure 6 demonstrates the performance variations across various influence function-based methods over 10 distinct attacks. Since the Logistic Regression model is convex, TDA simplifies to EKFAC, and TracIn and GEX simplify to IP; therefore, we will omit their duplicate results in this figure and the following table. In general, influence function-based methods are highly effective against adaptive evasion attackers, showcasing superior resilience particularly post-attack, notably evident in scenarios involving Bank and JigsawToxicity. Among these influence-based methods, our IP Ensemble demonstrates competitive efficacy with the best or the second best on all the datasets.
@@ 7 Defense Against Adaptive Evasion Adversaries
In addition to the trimming strategy, we continue to explore relabeling and reweighting strategies to tackle detrimental samples. The relabeling strategy [28] changes the identified detrimental samples from their original classes into another class, since the datasets we used here are binary classes, we directly flip their labels. The reweighting strategy [52] takes the influence score of each sample as the exponential weight with a softmax normalization, and then trains a model with weighted samples. We report the defense performance of various influence function-based methods and our IP Ensemble under the relabeling and reweighting strategies in Table 2. Under the relabeling strategy, DataInf achieves the best results on Bank and CelebA with 87.46% and 77.36% accuracy, respectively. However, LiSSA is not stable as other methods on CelebA with only 68.68%, which is even worse than the performance of post attack. Our IP ensemble method continues to achieve competitive performance, with accuracy scores of 87.45%, 77.30%, and 70.51% on three datasets with the second-best performance among all influence functions-based methods. For the reweighting strategy, the influence function-based methods are still effective on Bank and JigsawToxicity, which outperform the post-attack by a large margin, but achieve similar performance with post attack on CelebA. Notably, when considering the average accuracy of both the relabeling and reweighting strategies, our IP Ensemble method performs the best.
@@ 8 Conclusion
In this paper, we revisited, extended, and enhanced the TracIn method into the Inner Product (IP) formulation. By proving the exhibition of order-consistency in several scenarios between IP and influence function, we demonstrated substituting the inverse of the Hessian matrix with an identity matrix offers a practical and computationally efficient solution to estimating sample influence. Based on that, we extended our IP to measure the sample influence on fairness and robustness. Continually, we enhanced the generalization of IP by introducing an ensemble strategy. We verified the correctness of IP on synthetic datasets and extensive evaluations on noisy label detection, data curation for fair NLP model fine-tuning, and defense against adaptive adversarial attacks. Overall, our IP Ensemble highlighted the potential of simple, yet powerful, approximations in influence estimation.
@@ Appendix A Rationality of Inner Product
  • (I, I). In the scenario where both $z_{i}$ and $z_{i^{\prime}}$ have gradients in Region I, $\mathcal{I}^{\textrm{util}}(-z_{i})\gtrless\mathcal{I}^{\textrm{util}}(-z_{i^{\prime}})$ implies $\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i^{\prime}})\gtrless\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i})$, which is the opposite of order-consistency. While these samples are jointly recognized by $\mathcal{I}^{\textrm{util}}$ and $\mathcal{I}^{\textrm{util}}_{\textrm{IP}}$ as detrimental, there is little practical difference.
  • (I, II). In the scenario where $z_{i}$ has gradients in Region I and $z_{i^{\prime}}$ has gradients in Region II, the order-consistency partially holds, depending on which Sub-Region the gradients of $z_{i^{\prime}}$ is in. For Sub-Region $\mathrm{II^{b}}$ and $\mathrm{II^{c}}$, we have $\mathcal{I}^{\textrm{util}}(-z_{i})>\mathcal{I}^{\textrm{util}}(-z_{i^{\prime}})$ and $\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i})>\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i^{\prime}})$, and the order-consistency holds. For Sub-Region $\mathrm{II^{a}}$, however, $\mathcal{I}^{\textrm{util}}(-z_{i})\gtrless\mathcal{I}^{\textrm{util}}(-z_{i^{\prime}})$ and $\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i})>\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i^{\prime}})$, thus the order-consistency partially holds. While samples in Sub-Region $\mathrm{II^{a}}$ are jointly recognized by $\mathcal{I}^{\textrm{util}}$ and $\mathcal{I}^{\textrm{util}}_{\textrm{IP}}$ as detrimental, there is little practical difference.
  • (I, III). In the scenario where $z_{i}$ has gradients in Region I and $z_{i^{\prime}}$ has gradients in Region III, $\mathcal{I}^{\textrm{util}}(-z_{i})>\mathcal{I}^{\textrm{util}}(-z_{i^{\prime}})$ and $\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i})>\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i^{\prime}})$, therefore the order-consistency holds.
  • (II, IV). In the scenario where $z_{i}$ has gradients in Region II and $z_{i^{\prime}}$ displays gradients in Region IV, the concept of order-consistency is partially maintained.
  • (II, II). In the scenario where both $z_{i}$ and $z_{i^{\prime}}$ have gradients in Region II, $\mathcal{I}^{\textrm{util}}(-z_{i})\gtrless\mathcal{I}^{\textrm{util}}(-z_{i^{\prime}})$ implies $\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i})\gtrless\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i^{\prime}})$, the order-consistency holds.